Write the simplest polynomial function with the given roots: -3,2,4
Answer by ikleyn(52781)

.
It has the polynomial (binomial) factors (x-(-3)) = (x+3), (x-2) and (x-4).

The polynomial is this product p(x) = (x+3)*(x-2)*(x-4).

You may expand it, if you want/if you need.